# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

### v1

@@ 3.3 Active Endowment Generation
As established in Section 2, observed responses can serve as behavioral manifestations for latent preference vectors. Through this link, response variability serves as an initial diagnostic for assessing diversity in the underlying preference space. To effectively measure response variability, we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}$ (1)
where $H_{i}(\mathcal{A})$ denotes the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ is the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ is the empirical proportion of responses selecting option $k$. The denominator normalizes for scale, ensuring that entropy is comparable across questions with different numbers of choices. A high question entropy indicates greater response diversity for that question. By varying the group $\mathcal{A}$, e.g., restricting it to agents sampled from a particular mode, we can compute question entropy at the group level and compare the diversity contributes across modes.
To turn question entropy into a useful metric for adaptive sampling, we introduce variability score for each mode. Formally, it is defined as
$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}.$ (2)
where $H_{i}(\mathcal{A}_{\mathrm{mode}})$ denotes the entropy of responses to question $i$ based on the cumulative set of agents from a given mode up to the current generation step, and $H_{i}(\mathcal{A}_{\mathrm{total}})$ is the corresponding question entropy across all agents generated so far. Computed this way, the variability score favors modes that are capable of generating endowments which elicit diverse responses to questions that other modes typically fail to diversify. The intuition is to maximize the informational utility of each observation under budget constraints. When most agents agree on a question, a divergent response signals a valuable distinction in the underlying preference space. In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round. The scores are then passed through a softmax function to produce a probabilty distribution over modes, which is used to guide endowment sampling in the next generation step.
Adaptive sampling and update steps
In P2P, endowment generation takes several steps. At the outset of the generation, an attribute learner is invoked to infer attributes from the survey based on the training questions, yielding the survey template on top of the existing templates in the attribute bank. The system then enters an initial sampling stage, drawing an equal number of endowments from each mode. The endowments are then used to instantiate agents to elicit responses to training questions, $\mathcal{Q}_{\mathrm{train}}$. The collective responses are then passed to Tracker for the computation of the variability scores and the sampling probabilities for the next update step. At the next update step, endowment budget allocation follows the multinomial distribution. A higher probability exploits the top-performing mode, while randomness allows exploration of its under-performing peers.
@@
In addition to the existing trade-off between exploitation and exploration, we adopt an additional boosting strategy, which we call question patching. At the end of each update step, as the Tracker computes the variability scores for each existing mode, it also keeps track of the questions with the lowest entropy values. When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$n$ entropy questions, where $n$ is user-defined. The split of the endowment budget between the active generation and boosting generation is governed by a user-specified ratio.
Figure 3: Entropy change across update steps for each question during active endowment generation for ATP W42. Each column represents a question. Green indicates an increase in entropy from the previous step, while red indicates a decline.
To avoid wasting resources on persistently low-entropy and to promote cross-pollination between modes, we additionally introduce a mixed mode strategy. It is triggered after a question appears $n$ times in the lowest-$k$ entropy list, where $n$ is a user-defined threshold. When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode. This encourages diversity injection from high-performing modes while maintaining a question-specific focus.
The adaptive sampling cycle continues until the generation budget $N_{A}$ (an upper bound on the maximum number of agents allowed) is exhausted. By iteratively targeting low-diversity questions and expanding the attribute space, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space.
3.4 Regression-Based Aggregation
With the candidate proxy basis assembled, P2P enters the second stage: Regression-based Aggregation. The goal at this stage is to find a functional basis from the candidate vectors to reconstruct the observed preference. Depending on the endowment budget and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables than observations. Consequently, a variable selection algorithm is vital. Besides, even with $N_{\mathrm{train}}+N_{\mathrm{valid}}>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations. To conduct variable selection, we consider regression methods with an added L1 penalization term. We can use either hold-out validation or cross-validation (CV) for hyperparameter tuning. The objective of the learning problem can be stated as
@@ 3.4 Regression-Based Aggregation
where $\mathbf{w}\in\mathbb{R}^{N_{A}}$ is the weight vector for the agents, $r_{ik}$ denotes the ground-truth proportion of responses selecting option $k$ for question $i$, $d_{j,ik}$ is a dummy response variable equal to $1$ if agent $j$ selects option $k$ and 0 otherwise, and $\alpha$ is the hyperparameter for the penalization. When $\Omega$ consists of only a L1 norm, the formulation amounts to constrained lasso. When $\Omega$ subsumes an additional L2 norm, it becomes constrained elastic net. Our simulation studies indicates that constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations (Appendix A.1), and it is potent in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix A.2). Both results corroborate using it as a selection backbone. Our simulation study also investigates the relationship between entropy and prediction fit—we leave the details to Appendix A.3 for interested readers.
4 Taking P2P to the Wild
As a first empirical test of the alignment framework, we deploy P2P on real-world social survey data. Specifically, we use the American Trends Panel (ATP), a nationally representative U.S. survey conducted by Pew Research Center in discrete waves. In this demonstration, we choose Wave 42, which probes trust in science by asking respondents about scientific competence, integrity, and public accountability. The wave includes responses from approximately 4464 U.S. adults Pew Research Center (2019) and the cleaned dataset contains 128 questions.
For the experiment, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation and test sets. For the endowment generation stage, we set the endowment budget to be $300$, with $10$ endowments generated for each mode during initial sampling. Subsequently, the rest of the endowments are generated across 5 update steps. In each step, we allocate 75% of the endowment budget to question patching, where the lowest 3 questions are considered. We set the trigger threshold of the mixed mode strategy at 1, so that the budget is reallocated if a question enters the lowest-3 list the second time. At the end of the generation, 5 questions are patched, with additionally 5 mixed modes created.
Endowment generation improves entropy
Figure 3 displays the entropy change for each question across the update steps. Evidently, a significant number of questions experience a noticeable rise in entropy at the end of the first update step. From the second step onward, mixed mode enters the generation loop. Marginal gains in entropy are observed for the majority of questions, whereas some experience mild entropy drops. Appendix B provides question entropy trajectories in levels. The three questions that exhibit the largest drops in entropy are among the highest-entropy questions at the initial sampling stage. The evolution elucidates the tradeoff made by the active endowment generator encoded in the variability scores.
For the selection of the agent ensemble, we fit both a constrained lasso and a constrained elastic net, using CV to select the penalization parameters. Consequently, lasso selects 53 out of the 300 synthetic agents to form the agent ensemble, while elastic net selects 43. The training and testing performances are similar between the two methods.
Figure 4: Prompts to Proxies: Emulation Results for ATP W42. Top: Train (left) and test (right) prediction performances for the binarized responses. Bottom: Human and agent (pre- and post- alignment) aggregated responses for the sample test question RQ4_F1Ae_W42.
Prediction indicates feasibility
Figure 4 shows the prediction performances of the agent ensemble selected with constrained lasso. The train-time root mean squared error (RMSE) is 0.08, with test RMSE slightly higher at 0.09, indicating adequate generalization ability of the functional basis. The bottom panel of Figure 4 presents a snapshot for the aggregated responses for a test set question, comparing the human ground truth with agentic emulations before and after regression-based aggregation. Overall, the results are encouraging, especially considering that P2P constructs the endowments based solely on preset and learned attributes and is agnostic on the ground-truth demographic data.
The panoramic view and future directions
In Appendix B, we provide a preliminary panel study, applying P2P to 14 ATP waves, each repeated 3 times. Performance variation is relatively low within each wave but considerably higher across the panel. In 8 out of 14 waves, P2P consistently achieves a test MSE below 0.015 across all repeats. Four waves exhibit test MSEs fluctuating around 0.015, while the remaining two fall in the range of 0.020 to 0.025. While our simulation studies suggest that entropy coverage, the number of observations and the size of the proxy pool all influence emulation performance, the preliminary panel results signal intriguing directions for future study. Understanding the sources of variation in empirical performance—particularly in relation to both the similarities and the distinctions from simulation results—offers a promising lens on the dynamics of human–AI alignment. These insights can also inform refinements to P2P ’s modular architecture and motivate extensions of its functionality.
P2P as a means to an important end
At a broader level, our framework operationalizes pluralistic alignment Sorensen et al. (2024). Its structured use of attribute-based control enables the representation of diverse value systems, while its regression-based weighting mechanism supports distributional pluralism without enforcing demographic fidelity. Crucially, P2P demonstrates that steerability and pluralism are not separate challenges but deeply intertwined: the ability to steer model behavior through promptable attributes is what enables meaningful diversity in aggregate preference reconstruction.
Although developed in the context of survey emulation, P2P generalizes naturally beyond surveys. The underlying preference reconstruction logic can be applied to other forms of observational data by discretizing continuous variables and binarizing categorical ones. Moreover, while this paper focuses on modeling aggregate population preferences, the framework could be adapted for individual-level alignment—for example, by learning proxy ensembles that emulate preferences of a single person across varied contexts.
Beyond its methodological contributions, P2P affords a modular platform for empirical alignment. Our work is among the first to streamline the use of large-scale social survey data—not as textual corpora for language model pretraining, but as a structured, theory-informed substrate for steering models toward specific populations, cultural milieus, and value contexts. It supports systematic evaluation of prompt engineering techniques, attribute design strategies and sampling heuristics. Researchers can test how different endowment structures affect response diversity, and how entropy metrics reflect or obscure latent preference variation. These tools offer a bridge between high-level alignment goals and grounded behavioral evaluation.
Finally, P2P offers a testbed for social scientists to probe and operationalize preference theories. By investigating how assumptions about cognitive style, value orientation, or belief structure affect response distributions and mode-level entropy, researchers can explore the explanatory power to theoretical constructs in a controlled but flexible setting. In this sense, the system fulfills the spirit of its name: P2P connects peer disciplines, translates prospects into products, carries ideas from pseudocode to paper, and moves pluralistic alignment toward practical needs.

### v2

@@ 3.3 Stage 1: Active Endowment Generation
To measure preference diversity of agents in a revealed sense, we introduce the notion of question entropy:
$H_{i}(\mathcal{A})=\frac{-\sum_{k=1}^{K_{i}}p_{ik}\log_{2}p_{ik}}{\log_{2}K_{i}}\ ,$ (4)
where $H_{i}(\mathcal{A})$ denotes the normalized entropy of question $i$ based on responses from a group of agents $\mathcal{A}$, $K_{i}$ is the number of unique response options $\{1,\dots,K_{i}\}$ for question $i$, and $p_{ik}$ is the empirical proportion of responses selecting option $k$. The denominator normalizes for scale, ensuring that entropy is comparable across questions with different numbers of choices. A high question entropy indicates greater response diversity for that specific question. By restricting the group $\mathcal{A}$ to agents sampled from a particular mode (subject area), we can compute question entropy at the mode level, termed mode-question entropy, and compare the response variability across modes.
Adaptive sampling references the variability score for each mode. Formally, it is defined as
$V(\mathcal{A}_{\mathrm{mode}})=\sum_{i\in\mathcal{Q}_{\mathrm{train}}}\frac{H_{i}(\mathcal{A}_{\mathrm{mode}})}{H_{i}(\mathcal{A}_{\mathrm{total}})}\ ,$ (5)
where $H_{i}(\mathcal{A}_{\mathrm{mode}})$ denotes the entropy of responses to question $i$ based on the cumulative set of agents from a given mode up to the current generation step, and $H_{i}(\mathcal{A}_{\mathrm{total}})$ is the corresponding question entropy across all agents generated so far. Computed this way, the variability score favors modes that are capable of generating endowments which elicit diverse responses to questions that other modes typically fail to diversify. In P2P, we task a dedicated class object, Tracker, to compute the variability score for each mode after each generation round (see Algorithm 2 in Appendix B.3). The scores are then passed to a softmax function to produce a probabilty distribution over modes, guiding endowment sampling in the next generation step via a multimonial draw—balancing exploitation of top-performing modes with exploration of under-performing ones.
@@
In addition to the existing trade-off between exploitation and exploration, we adopt an additional patching strategy, which we call question patching. At the end of each update step, as the Tracker computes the variability scores for each existing mode, it also keeps track of the questions with the lowest entropy values. When question patching is activated, the attribute learner will be called in the next generation step to exclusively infer attributes on each of the lowest-$k$ entropy questions, where $k$ is user-defined. The split of the endowment budget between active generation and patching generation is governed by a user-specified ratio.
To avoid wasting resources on persistently low-entropy and to promote cross-pollination between modes, we additionally introduce a mixed mode strategy. It is triggered after a question appears $t$ times in the lowest-$k$ entropy list, where $t$ is a user-defined threshold. When in effect, the strategy relocates the assigned endowment budge to a new mixed mode, which combines attributes from the question with those of the current top-performing mode, allowing diversity injection from high-performing modes while maintaining a question-specific focus. Figure 10 shows example endowments generated from a real run of P2P, illustrating how different templates and modes yield distinct profiles.
The adaptive sampling cycle continues until the generation budget $N_{A}$ (an upper bound on the maximum number of agents allowed) is exhausted. By iteratively targeting low-diversity questions, expanding the attribute space and promoting cross-pollination between modes, active endowment generation progressively increases response diversity while preserving coverage over previously explored regions of the preference space (cf. Figure 5 in Appendix B.4).
3.4 Stage 2: Regression-Based Aggregation
With the proxy basis assembled, P2P enters the second stage: regression-based aggregation. The goal at this stage is to find a functional basis from the proxy basis to reconstruct the population preference based on the observational data. Depending on the endowment budget $N_{A}$ and the total number of trainable questions $\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}$, we could end up in a situation where there are more variables (proxy agents) than observations (trainval questions). Consequently, a variable selection algorithm is vital. Besides, even with $|\mathcal{Q}_{\mathrm{train}}\cup Q_{\mathrm{valid}}|>>N_{A}$, variable selection allows us to retain a parsimonious agent ensemble, saving inference costs on further simulations. We consider regression methods with an added L1 penalization term and use cross-validation (CV) for hyperparameter tuning. The objective of the learning problem can be stated as
@@ 3.4 Stage 2: Regression-Based Aggregation
where $\mathbf{w}\in\mathbb{R}^{N_{A}}$ is the weight vector for the agents, $r_{ik}$ denotes the ground-truth proportion of responses selecting option $k$ for question $i$, $d_{j,ik}$ is a dummy response variable equal to $1$ if agent $j$ selects option $k$ and 0 otherwise, and $\alpha$ is the hyperparameter for the penalization. When $\Omega$ consists of only a L1 norm, the formulation amounts to constrained lasso. When $\Omega$ subsumes an additional L2 norm, it becomes constrained elastic net.
To validate L1-regularized regression as a credible aggregation algorithm, we conduct extensive simulation studies for constrained lasso (Appendix C), where ground truth population preference is constructed using synthetic agents [^9] This allows us to create a clean environment where alignment result is not affected by the learnability condition (Def. 2.2)..
Our simulation studies yield three main findings. First, constrained lasso faithfully recovers the ground-truth agents from a large candidate pool under an adequate number of observations, including minority profiles with low weights, and only resorts to reconstruction when observations are extremely scarce (Appendix C.1). Second, when ground truth agents are removed from the candidate pool, constrained lasso remains effective in selecting a proxy ensemble to recover ground truth even when the number of proxy agents is small (Appendix C.2). Third, proxy ensembles with insufficient response entropy fail to recover ground-truth preferences, while those with excess entropy show only marginal deterioration, corroborating our choice of response variability as the optimization goal of Stage 1 of P2P (Appendix C.3). Together, these results provide constructive evidence for P2P’s design choices.
4 Empirical Evaluation
Having validated P2P’s design choices through simulation, we now evaluate it on real human survey data. We use the American Trends Panel (ATP) (Pew Research Center, ), a nationally representative U.S. survey conducted by Pew Research Center. We obtain the digitalized versions of these datasets from the OpinionsQA benchmark (Santurkar et al., 2023), which repurposed ATP responses to study value pluralism and misalignment in language models. We assess P2P across 14 ATP waves, spanning diverse socio-economic topics. We then evaluate locale generalizability on the World Values Survey, conduct error analysis and ablation studies, and stress-test against a distribution-calibrated model under topic shift.
4.1 Experimental Setup
For each survey, we split the questions according to a $7:1.5:1.5$ ratio into the training, validation, and test sets. To benchmark P2P, we create two prompting-based ensemble baselines with different endowment generation logic but feed both through the same regression-based aggregation stage as P2P for proxy selection and test set evaluation. This allows the baselines to benefit from the same data-driven selection mechanism, isolating the contribution of P2P’s endowment generation strategy.
Vanilla baseline
Endowments are generated using an LLM prompted [^10] The endowment generation prompts for P2P and the Vanilla baseline can be found in Appendix B.2. directly to generate diverse agent profiles conditioned only on the survey topic.
PERSONA
Endowments are sampled randomly [^11] We use random sampling as it preserves the representativeness of the dataset. In Appendix D.1.1, we provide baseline variants using K-means clustering as a pre-processing step. Clustering does not improve performance but increases cost. from the demographically conditioned PERSONA dataset (Castricato et al., 2025), which contains 1000 persona profiles derived from U.S. census data.
For all methods, we set the endowment generation budget to $300$. For P2P, initial sampling generates $10$ endowments for each mode; active endowment generation is run in 5 update steps with the preset attribute bank, question patching (lowest 3; 75% endowment budget), and 10 attributes drawn for each endowment generation. Unless otherwise noted, all models use Gemini-2.0-flash as the backend [^12] See Appendix D.1.2 for raw model performance., and constrained lasso is used for Stage 2 selection.
4.2 Main Results: Baseline Comparison on ATP Waves
Table 1: Performance comparison across 14 ATP waves. Test MSE, entropy, and cost (USD) reported. P2P reports mean over 3 runs. Baselines report single runs, as repeated runs on W27, W32, W42, and W45 show negligible variance (MSE std << 0.005), confirming stability (Appendix D). $\Delta$ shows relative test MSE improvement over the best baseline.
Vanilla PERSONA P2P
Wave Topic # Qs MSE Ent. Cost MSE Ent. Cost* MSE Ent. Cost $\Delta$
W26 Guns 78 .015 .50 \$0.5 .021 .37 \$1.0 $\mathbf{.010}$ .55 \$0.6 33%
W27 Auto 96 .030 .40 \$0.6 .042 .40 \$1.3 $\mathbf{.022}$ .54 \$0.7 27%
W29 Gender 77 .030 .45 \$0.5 .030 .44 \$1.0 $\mathbf{.016}$ .51 \$0.6 47%
W32 Comm. 98 .020 .55 \$0.6 .026 .56 \$1.3 $\mathbf{.016}$ .62 \$0.7 20%
W34 Biomed 67 .022 .43 \$0.5 .020 .39 \$0.9 $\mathbf{.011}$ .54 \$0.5 45%
W36 Lead. 139 .033 .34 \$0.9 .035 .39 \$1.8 $\mathbf{.019}$ .48 \$1.0 42%
W41 2050 90 .022 .45 \$0.6 .023 .48 \$1.2 $\mathbf{.013}$ .55 \$0.8 41%
W42 Science 129 .034 .32 \$0.9 .038 .29 \$1.7 $\mathbf{.011}$ .43 \$1.0 68%
W45 Misinfo 95 .016 .39 \$0.6 .028 .36 \$1.2 $\mathbf{.008}$ .44 \$0.7 50%
W49 Privacy 98 .045 .41 \$0.6 .047 .41 \$1.3 $\mathbf{.021}$ .53 \$0.7 53%
W50 Family 128 .027 .36 \$0.8 .024 .36 \$1.7 $\mathbf{.013}$ .44 \$0.9 46%
W54 Econ 116 .022 .50 \$0.8 .026 .54 \$1.5 $\mathbf{.014}$ .57 \$0.9 36%
W82 Global 104 .025 .52 \$0.7 .019 .48 \$1.4 $\mathbf{.008}$ .57 \$0.8 58%
W92 Poli. 77 .025 .59 \$0.5 .031 .57 \$1.0 $\mathbf{.017}$ .61 \$0.6 32%
Average .026 .45 \$0.7 .029 .43 \$1.3 $\mathbf{.014}$ .53 \$0.8 43%
* PERSONA cost only covers the response elicitation stage—higher if persona generation tokens are included.
As Table 1 shows, P2P consistently outperforms both baselines, achieving an average test MSE of 0.014 compared to 0.026 (Vanilla) and 0.029 (PERSONA)—a 43% improvement over the best baseline. For P2P, 12 out of 14 waves achieve mean test MSE (across 3 runs) below $0.02$, confirming generalization across topics without wave-specific tuning. The agent pool created using P2P also achieves higher diversity than the baselines, indicated by the higher average entropy (0.53 vs. 0.43-0.45), corroborating our simulation finding that response diversity supports preference reconstruction (Appendix C.3). Notably, this diversity gain is cost-efficient, as P2P runs at approximately \$0.8 per survey, similar to Vanilla (\$0.7) and substantially cheaper than PERSONA (\$1.3, excluding persona generation overhead) [^13] The PERSONA pipeline (Castricato et al., 2025) requires resampling and re-inference, adding further API usage costs.. Performance comparison is robust under other distributional metrics (1-JSD, EMD) (Appendix D).
4.3 Locale Generalization: World Values Survey
To probe whether P2P generalizes beyound U.S. populations and ATP-specific survey design, we evaluate it on World Values Survey (WVS) Wave 7 (Haerpfer et al., 2022), using cleaned data from Cao et al. (2025). We focus on three representative regions: the United States, Great Britain, and Hong Kong SAR. Under the same experimental protocol [^14] We pass the locale information to the endowment model to ensure regional relevance of the generated endowments., P2P achieves comparable performance on US and GB (MSE below $0.015$), with somewhat weaker results on HK (MSE around $0.024$) (Table 2).
Table 2: P2P performance on WVS Wave 7 across three locales. Values are mean $\pm$ std over 3 runs. (Full details in Appendix D.)
Metric US GB HK
Entropy $\uparrow$ .572 $\pm$ .015 .571 $\pm$ .004 .548 $\pm$ .021
Test MSE $\downarrow$ .015 $\pm$ .001 .014 $\pm$ .001 .024 $\pm$ .001
Cost $\downarrow$ 1.76 $\pm$ .01 1.71 $\pm$ .00 1.74 $\pm$ .01
Figure 3: Error Analysis of Hong Kong WVS Results. We compare UK-native and HK-native with HK-transfer (US endowments applied to HK via P2P Stage 2). (A) Ground truth divergence between populations ($r=0.848$). (B) Transfer outperforms native in overall performance, corroborating the functional basis perspective. (C) Error gap scales with cultural divergence, suggesting learnability limits (Definition 2.2). (D) Despite higher test MSE, HK-native outperforms transfer on culturally divergent questions, indicating that localization helps where cultures genuinely differ.
4.4 Error Analysis: Transfer vs. Native Alignment
One plausible explanation for the WVS performance gap is that LLMs trained primarily on English and Western-centric data better reflect Western populations (Durmus et al., 2024). Yet, we want to probe it further by connecting to our theoretical framework: How much of the gap is explained by insufficient agent diversity and can we characterize the learnability limits? To address these questions, we take US endowments, which have higher entropy, and apply them to HK via P2P Stage 2 by refitting regression weights. We call this HK-transfer and compare it against those in Table 2 (called US-native and HK-native).
Figure 3 presents four complementary analyses. We define cultural divergence for each question-option pair as $|\mathrm{GT}_{\mathrm{US}}-\mathrm{GT}_{\mathrm{HK}}|$, the absolute difference in response proportions between populations. Panel (A) shows correlated but distinct US and HK ground-truth response patterns ($r=0.848$), with variation around the diagonal confirming cultural differences. Panel (B) shows that HK-transfer achieves lower test MSE than HK-native, corroborating the functional basis perspective—a more diverse agent pool facilitates preference reconstruction, even when generated for a different locale [^15] This result cuts against demographic conditioning approaches, which assume locale-matched profiles are necessary for faithful preference reconstruction.. However, Panels (C) and (D) reveal important nuances. Panel (C) shows that prediction error scales with cultural divergence, offering an empirical handle on learnability limits (Definition 2.2). In culturally distinct dimensions, Gemini-based ensembles exhibit a clear bias toward US ground truth. As indicated by the trends, conditioning on locale information provides a mild mitigation. Panel (D) further corroborates this, showing that despite higher test MSE, HK-native outperforms HK-transfer on culturally divergent questions. The takeaway is that while cross-locale transfer can leverage diversity on average, locale-specific generation may capture culturally distinctive preferences that a foreign basis cannot fully reconstruct. The analyses also demonstrate P2P’s potential as a benchmark for model steerability and bias research (cf. Appendix F).
4.5 Ablation Studies
We conduct ablation studies on ATP Wave 42 to isolate component contributions. Full results are in Appendix D.3.
Attributes
Disabling question patching reduces entropy from 0.445 to 0.399 and increases test MSE from 0.010 to 0.011, while disabling both freeform templates (question & survey) further degrades MSE to 0.013. This suggests that human-AI co-discovery of preference descriptors yields meaningful performance gains. Varying the per-endowment attribute cap (10 vs. 20) has minimal effect.
Endowment budget
Performance improves with budget (110-450 agents) but plateaus after 3 update steps (210 agents). A no-update baseline (110 agents) underperforms, validating the benefit of adaptive sampling.
Model backend
We test P2P’s performance on five model backends: GPT-4.1-mini, GPT-4.1-nano, Gemini-2.0-flash, Gemini-2.5-flash and a locally hosted Qwen. GPT-4.1-mini achieves the best entropy (0.471) and test MSE (0.009). Gemini-2.0-flash ranks second at one-quarter the cost. Smaller models (GPT-4.1-nano and Gemini-2.5-flash-lite) perform worse, consistent with our intuition that limited model capacity constrains the learnability (Definition 2.2).
Regression-based aggregation
Replacing regression with simple averaging of agent responses increases test MSE from 0.0104 to 0.0254, confirming that regression is essential for learning population preferences, not just ensemble compression. Switching loss weighting from option level to question level has minimal effect on W42 results.
4.6 Stress Test: Comparison with an SFT-Aligned Model under Topic Shift
P2P is designed as an inference-based alignment method for limited-data and limited-compute regimes, so our baselines focus on prompting-based methods under similar constraints. However, supervised finetuning (SFT) offers an alternative when more data is available. For instance, Cao et al. (2025) train a distribution-calibrated model on pooled WVS data from 46 countries (approx. 8,400 training entries) to predict answer distributions conditioned on locale.
We stress-test P2P on their protocol, which induces a topic shift: training on general attitudinal and religious-ethical items (Q1+Q2) and testing on political-cultural items (Q3). We apply P2P to three locales (the United States, Great Britain, and Hong Kong SAR), using only 172-185 local training entries and no finetuning. Despite using less than 3% of the training data and no SFT, P2P achieves comparable Jensen-Shannon Divergence (1-JSD) (0.71-0.75 vs. 0.78), with matching or better majority-class prediction accuracy (MCPA) (0.47-0.64 vs. 0.43). We interpret this as evidence that some preference structure transfers across topics though we do not assume such transfer is guaranteed. Studying cross-topic generalization remains an important direction for future work (Appendix F). Details on the stress test are presented in Appendix D.4.